OpenStudy (anonymous):

Find the domain of the function and express the answer in interval notation: g(x)=(2x)/(x-3)

OpenStudy (anonymous):

i would say that the answer is (-\[\infty\], 3]

OpenStudy (anonymous):

When x=3, g(x) is undefined. Therefore the domain is \[(-\infty,3)\cup(3,\infty)\]

OpenStudy (anonymous):

ok, so it can be any reall number expect 3?